Dan acquired rental property in June 2008 for $370,000 and sold it in October, 2018. $40,000 in straight-line depreciation has been taken on the house. A run-up in housing prices allowed him to sell the house for $575,000. In the year of sale, Dan received $175,000 when the buyer sold some investments, an additional $200,000 when the buyer closed a loan from the bank, and took a $200,000 note from the buyer, payable on the anniversary of the sale date in 10 installments of $20,000 each plus interest on the unpaid balance.

Using the installment method, calculate his taxable gain in the year of sale. **For year 2018**
Answer #1

In order to solve this example we need to understand firstly the meaning of installment method, lets see below:

  • Installment method which in the above case is installment sale is a type of sales agreement wherein the seller gives the benefit to the buyer to postpone the payment for the actual sale transaction.
  • The main purpose of such an agreement is that the seller can defer his long term capital gain to the future taxation years and will also receive regular installment amounts along with the benefit of the interest on the unpaid capital balance.
  • For the buyer it gives him immediate ownership of the asset/property along with sufficient time to pay back his debt to the seller

In our case there is a note payable payable on the anniversary of the sale date in 10 installments of $20,000 each plus interest on the unpaid balance, which means that installment payment for the note will begin from Jan 2019 and will be pain in equal 10 installments till the anniversary sale date (in our case sale date is Oct 2018 so the anniversary will finish in Oct 2019) , hence if you see here Dan would have not received any amount of installment or interest on installment for the year 2018 as the installment payment starts from Jan 2019.

So now lets see what will be his taxable gain:

Cost of rental property as on June 2018 - $ 370,000

(-) Depreciation cost which is mentioned -$ 40,000(the rental property is held with Dan from June 2008 to Oct 2018 which is 10 years so the amount charged against depreciation each year would have been 40000/10=4000 per year as per straight line method

Net cost of the rental property will be 370,000-40,000 = 3,30,000

However the property is sold in Oct 2018 at $ 575,000 so the net gain from the sale of rental property is (575000-330000(cost minus depreciation) = 2,45,000

we will be computing the taxable gain on $245,000 up to the extent of the cash receipts received by Dan for the year 2018 and he only receives a total of 375,000 from the buyer (sale of investment 1,75.000 and closing of loan 2,00,000) and the balance which he will receive from installment sale of 2,00,000 will not be received in the year 2018 so lets compute the ratios:

375000/575000=0.65

Taxable gain for Dan for the year 2018 = 2,45,000 * 0.65*20% = 31,850

Here the long term capital gain on the rental property being 20% we have charged total percent above.

Your feedback is valuable to us, please give a thumbs up if you feel so.